Workers: Miyo, Akutsu
- Restarted the cryocon.
- Yokozawa-san reported it was strange on Sunday.
In short, restarting it did not help the situation very well. No idea...
- Anyway EPICS can pick up the data from it. So far so ok.
- Attached the sixth, and the last thermometer at the side pillar of the suspension.
- The place is opened when the protectors were removed. So installed today.
- The signal wires were connected to the cryocon, and now online.
- Note that this thermometer uses all 75-ohm lines, different from the other ones.
- Heat-anchored some electric wires to the IYC chamber.
- One is the wires from the thermometer described above; Kimura-san said we don't need to care, but I can do so I did.
- Another is fluffy wires; to fix them I just press.
[Ongoing] Need cares of water springs from anchor bolts for oplev pylons.
[Ongoing] Need to confirm what are the scratch lines made by the measurement by the KEK team last week.
- Make heat connection upper and lower part of the suspension by Cu wires.
- Complete the PS calibration.
- Complete the calibration of the oplev; the arm length was reported by Simon, so....
- Complete the signal matrix.
- Check the balance of the suspended WAB with some wires and sensors are attached.
- Check how the damping would be now after the replacement of the damping magnets.
- Check the reproducibility of the alignment of the suspension among several modes.
- Check heights of several points.
- Take care of water spring??
- Two burndy connectors, vacuum compatible.
Stages outside the chamber to ease wiping our legs.
Good rail(?) to move the WAB suspension in the cryostat in the main beam direction.
Chain block? to support the load in the chamber.
- A clock to know the time easily.
- M6 screw set.